Find two consecutive even integers such that 3 times the smaller is 12 more than 2 times the larger.
Answer by checkley77(12569)

Let x & (x+2) be the 2 integers.
3x=2(x+2)+12
3x=2x+4+12
3x-2x=16
x=16 answer for the smaller integer.
16+2 for the larger integer.
Proof:
3*16=2*18+12
48=36=12
48=48